OpenStudy (anonymous):

CD is tangent to the circle and segment AE is perpendicular to segment AC. If a radius of the circle measures 12 in., which is the vale of DF + AF?

OpenStudy (anonymous):

Sorry but im terrible at math

OpenStudy (turingtest):

these must be 3:4:5 triangles (30:60:90 degrees) so AE=R=12 AF=(3/4)AE=9 CF=R-AF=12-9=3 DF=(5/3)CF=5 AF+DF=14 I'm sorry I didn't show my work very clearly, I got a little distracted.